OpenStudy (anonymous):

How do I eliminate x from the following system? a*tan(x) +1 = sec(x) b*tan(x) -1 = sec(x)

OpenStudy (anonymous):

Why would you want to do that?

OpenStudy (anonymous):

come on that's your answer?

OpenStudy (anonymous):

Eliminate it and leave what?

OpenStudy (anonymous):

the answer is a*b=1 but i dont know how to get that

OpenStudy (anonymous):

You have a = secx-1/tanx and b = sec x+ 1/tan x so ab = (sec x-1)(secx+1)/tan^2x = 1

OpenStudy (anonymous):

Thank you, I didn't try to just let the a and b alone.